### v2

@@ 3.1 Limitations of Existing Data Resources
Moreover, our analysis reveals that existing resources provide highly unbalanced coverage of various cultural aspects, with a notable lack of data on social norms and commonsense, as shown in Figure 3. We manually classify all questions into one of five categories: (1) Cultural Entities, where the question asks about culture-specific entities Naous and Xu (2025), (2) Cultural Opinion, where the question asks about a subjective interpretation for a cultural entity, (3) Social Norms, where the question is about social interactions between individuals Huang and Yang (2023), (4) Cultural Commonsense, where the question is about daily phenomena that locals may take for granted Shen et al. (2024), and (5) Literacy, where the question is about language usage, proverbs, or slang Wuraola et al. (2024). Figure 2 shows examples for these cultural aspects. Inter-annotator agreement is 0.90, measured by Cohen’s Kappa, for this classification.
Finally, aligning LMs using preference optimization algorithms requires human judgments on both suitable and unsuitable responses, beyond only the existing ground-truth demonstrations.
3.2 Multilingual Cultural Alignment Data
To address the aforementioned shortcomings, we collect new non-English data for cultural alignment as follows:
Social Norm and Commonsense.
To expand the samples on social norms and cultural commonsense, we ask native Chinese and Arabic speakers who are international college students undergoing culture transfer experiences themselves to curate such samples manually. To help brainstorm, the speakers are instructed to leverage international as well as regional social media platforms or forums (e.g., Twitter, Reddit, Zhihu, RedNote), and search for posts where users describe their culture shift experiences using search keywords such as “most surprised abroad, culture shock, first time to the US, etc.”. Taking inspiration from such discussions, the speakers create 190 samples in Chinese and 196 samples in Arabic. This human-curated data is more authentic than synthetic data, preventing possible inaccuracy and overgeneralization (e.g. “In China, it is common to drink soup after the main dish.”) we have observed in the existing datasets.
Culture-specific Human Preference.
Given these culture-specific questions, gathering human preferences on the cultural relevance of responses is crucial to performing cultural alignment of LMs through preference optimization. To obtain these cultural preferences, we present the native annotators with the zero-shot responses of 9 different LMs for each question within CARE. We specifically use the instruct version of recent multilingual LMs of Llama3.3-70B Dubey et al. (2024), Qwen2.5-72B Yang et al. (2024), Gemma2-27B Team et al. (2024), Mistral-Large, and GPT-4o. We also use their smaller-sized Llama3.1-8B, Qwen2.5-7B, Gemma2-9B, and Mistral-7B.
We then instruct the annotators to rank the generated responses for each question from the most to least culturally appropriate and assign each response a rating on a scale of 1 to 10 (1: poor $\rightarrow$ 10: excellent). Those ratings reflect how well responses match the cultural expectations of native speakers, which we use to construct preference pairs for cultural alignment in our experiments (§4.1). For responses that deviate from answering the questions (e.g., generating similar questions or repetitive content), annotators do not assign a rating.
Human Annotation.
We hire two native Chinese speakers and two native Arabic speakers who are college students to collect and annotate all the data in CARE, as described in §3.1 and §3.2. Besides the 5-class cultural categories, we also annotate each sample with the Associated Culture for experiments in §4.5: Native (questions about the native culture; i.e. Chinese or Arab), Foreign (questions about other cultures), or General (questions that are not specific to a particular culture). For further fine-grained regional cultural knowledge evaluation (Appendix D.1), we also annotate samples for their Geographic Scope (sub-nationwide, nationwide, continent-wide, or worldwide). See Appendix A.1 for the annotation guideline.
Data Split.
For the experiments in §4 and §5, we construct culture-specific test sets with 150 questions each for Chinese and Arab cultures, randomly sampling 30 questions from each of the five cultural categories. The remaining 1,513 Chinese and 644 Arabic questions, along with their responses in CARE, are used to form the training set. This training data also includes questions about other foreign cultures beyond Chinese and Arab.
Chinese Arabic Model Entities Opinion Norms C. sense Literacy Average Entities Opinion Norms C. sense Literacy Average [Uncaptioned image] Gemma2-9B Vanilla 4.89 8.46 7.55 6.67 4.72 6.49 4.73 6.56 6.33 5.60 3.28 5.33 Aligned[Uncaptioned image] 5.50 8.53 7.90 7.26 5.17 6.89 5.30 6.82 6.93 6.55 3.57 5.84 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Llama3.1-8B Vanilla 3.14 4.16 4.62 3.93 3.03 3.78 4.08 3.62 2.87 3.07 2.07 3.30 Aligned[Uncaptioned image] 3.86 5.90 5.36 5.56 3.69 4.88 4.33 4.43 3.70 4.50 2.36 3.86 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Qwen2.5-7B Vanilla 6.89 7.86 7.48 6.80 7.37 7.28 4.65 5.84 5.44 4.88 2.84 4.61 Aligned[Uncaptioned image] 7.20 8.76 7.66 6.90 7.53 7.61 4.55 6.40 5.55 5.33 3.35 5.06 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Mistral-7B Vanilla 3.03 3.83 3.93 4.38 2.43 3.53 2.56 2.46 2.03 2.13 1.34 2.11 Aligned[Uncaptioned image] 2.43 3.90 4.53 5.00 2.20 3.61 2.60 3.36 2.46 2.10 1.40 2.38 [Uncaptioned image] [Uncaptioned image]
Table 1: Average scores (1: poor $\rightarrow$ 10: excellent) in responding to questions related to Chinese culture in Chinese and Arab culture in Arabic. Performances are presented for vanilla LMs and LMs after cultural alignment using DPO on CARE[Uncaptioned image]. For each LM, the row labeled “Vanilla” corresponds to the original model (gray plot), and “Aligned” is after cultural alignment (colored plot). “C. sense” stands for “Commonsense”.
@@ 3.2 Multilingual Cultural Alignment Data
4 Aligning LMs for Cultural Awareness
Enabled by CARE, we investigate whether a small amount of high-quality cultural alignment data can improve existing models through preference optimization. We first conduct cultural alignment on medium-sized LMs and show how much improvement we can achieve with CARE in §4.1 and §4.2. Next, we examine the impact of scaling data and model size (§4.3) and whether cultural alignment sacrifices LMs’ capabilities on general tasks (§4.4). Finally, we analyze the importance of including data from diverse cultures during alignment (§4.5).
4.1 Experiment Setup
We perform human preference learning for cultural alignment of medium-sized LMs (7$\sim$9B parameters) using CARE on up to 8 Nvidia A40 GPUs.
Human Preference Optimization.
We experiment with different preference learning algorithms, including DPO Rafailov et al. (2024), KTO Ethayarajh et al. (2024), and SimPO Meng et al. (2024). These algorithms require preference pairs that contrast two responses for each question: one culturally appropriate and one that is less appropriate. We use the highest- and lowest-scored responses from the human preference data collected in CARE. We find that all three offline preference algorithms perform similarly well, as shown in Table 3.2; subsequently, we present most of the experiment results in this paper with DPO unless otherwise specified.
Baselines.
First, we analyze the effectiveness of learning from cultural preference compared to general preference. We align LMs using existing human preference datasets, such as OpenOrca Lian et al. (2023) and UltraChat Bartolome et al. (2023), which focus on truthfulness, honesty, and helpfulness. For a fair comparison, we randomly sample an equal number of training pairs from these datasets as in CARE’s training set. Second, we examine the effectiveness of SFT on CARE versus Alpaca Cui et al. (2023); Chen et al. (2023), a general instruction-tuning dataset with Chinese and Arabic adaptations. Third, for prompting-based approaches, we explore CoT prompting Wei et al. (2022) and role-play prompting Kong et al. (2023). See prompts used in Appendix B.
LM-as-a-judge Evaluation and its Reliability.
For evaluation, we adopt the LM-as-a-judge Zheng et al. (2023) strategy where we prompt a powerful LM to evaluate how aligned the generated response is to the human reference response. We use GPT-4o as the LM judge and prompt it to provide reasoning and a score on a scale of 1 to 10. We validate this setup by comparing it to the ratings of the native-speaking human annotators, where we achieve a high Pearson correlation of 0.93 between human and GPT-4o ratings. More details and our evaluation prompts can be found in Appendix C.
[…]
Table 3: Average scores (1: poor $\rightarrow$ 10: excellent) on Chinese and Arab cultures for a variety of prompting approaches, fine-tuning, and alignment using cultural (CARE) vs. generic instruction-tuning (Multilingual Alpaca) and preference (OpenOrca/UltraChat) data. SFT is performed on the instruction data only, while alignment conducts DPO on the human preference pairs. Aligning with cultural preferences enhances cultural awareness more effectively than using general preferences, with this improvement remaining consistent across various prompting setups.
@@ 4.2 Main Results
Table 1 presents the results of medium-sized LMs before and after cultural alignment using CARE, while Table 3 compares these results with various baselines. We have the following key findings:
Using CARE, culturally-aligned LMs achieve higher average scores (up to 29% improvement) in both Chinese and Arabic compared to the vanilla checkpoints, even for these already instruction-tuned models. This resonates with the findings of Zhou et al. (2024a) and shows that a relatively small amount of carefully curated data by humans can improve LMs’ alignment. We also see a noticeable gap among LMs developed by different regions. The Qwen2.5-7B developed by China-based Alibaba performs the best in Chinese and can be further improved from 7.28 to 7.61 (out of 10) overall by aligning with CARE. Interestingly, the aligned version of Gemma2-9B outperforms Qwen2.5-7B on social norms and commonsense for Chinese (more on this in §5.2). We also test its generalization capabilities on other culture-related tasks in Appendix D.
Cultural alignment strengthens stronger LMs but fails to address weaknesses in weaker LMs. Qwen2.5-7B and Gemma2-9B, the top-performing LMs on Chinese and Arabic, respectively, show consistent improvement across all five cultural categories after alignment. However, alignment shows limited gains where the model’s initial performance is poor. Mistral-7B is not improved on Chinese entities and literacy, where its starting scores are only 3.03 and 2.43. Similarly, Qwen2.5-7B, does not benefit from alignment on Arabic entities. This suggests that base models need a foundational level of cultural knowledge for alignment to be effective.
Culture-specific human preference data is helpful beyond general human preference. Table 3 shows that alignment with CARE’s cultural preference consistently outperforms general preference datasets by up to 42%. Additionally, cultural alignment with CARE brings improvement in both CoT and role-play prompting. While culture-specific SFT outperforms generic instruction tuning, alignment yields greater gains, highlighting the importance of human cultural preference data over ground-truth demonstrations alone.
4.3 How does scaling affect alignment?
We examine how scaling both model size and data volume influences the performance before and after cultural alignment. Specifically, we consider the Qwen2.5 {1.5B, 3B, 7B, 14B} and the Gemma2 {2B, 9B} series, which offer multiple model sizes. For data sizes, we align models using DPO on different proportions {0%, 25%, 50%, 75%, 100%} of Chinese cultural preference pairs within CARE.
As shown in Figure 4, consistent improvements are observed across different model and data sizes. Even a relatively small amount (e.g., 25%) of data in CARE can lead to improvements, particularly for smaller-sized LMs. Scaling up the data leads to progressively better performance, highlighting the potential benefits of employing more cultural preference data in future work.
4.4 Does alignment impact general capability?
We then examine whether cultural alignment impacts the model’s overall knowledge and capabilities, using well-established benchmarks for Chinese, Arabic, and English: ArabicMMLU Koto et al. (2024), ChineseMMLU Li et al. (2023), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2021), and WinoGender Rudinger et al. (2018). We compare different versions of Llama3.1-8B-Instruct: the vanilla LM, the aligned LM with general human preference from the combined UltraChat and OpenOrca datasets (totaling 3,718 samples), the culturally aligned LM with cultural human preference from CARE, and the aligned LM with both general and CARE’s cultural preference.
The results in Figure 5 show very small differences between the vanilla LM and its aligned version, on benchmarks in both native language and English, suggesting that cultural alignment does not hinder the model’s overall capabilities.
Figure 5: Accuracy of Llama3.1-8B-Instruct variants on general benchmarks assessing overall capabilities in native languages (Chinese and Arabic) and English. Slight differences are observed among the LM variants after applying different alignment setups.
4.5 Should we align with only native data?
We analyze the impact of incorporating data about different cultures in the alignment process. To ensure a fair comparison, we select an equal number of samples for each trial. Specifically, we align Llama3.1-8B-Instruct with 676 Chinese and 546 Arabic pairs sampled from three different contexts in CARE: native (Chinese or Arab) culture, foreign cultures (non-Chinese or non-Arab), and mixed pairs (half native and half foreign).
Results are shown in Figure 6. We find that performance drops on native culture when the model is aligned using only foreign culture samples, but improves when aligned with the same native culture samples (4.69 $\rightarrow$ 4.87 in Chinese and 2.98 $\rightarrow$ 3.50 in Arabic); interestingly, performance is further improved when the model is aligned by combining native and foreign cultural data. This indicates that expanding geographical coverage in alignment could further enhance cultural awareness.
[…]
Table 4: Average scores on CARE of larger LMs and web content, evaluated by the judge LM across all samples (1: poor $\rightarrow$ 10: excellent). “Human Awareness (%)” indicates the percentage of questions for which natives know the correct answer. Qwen2.5 developed by Chinese-based Alibaba excels on Chinese entities, norms, and literacy, while GPT-4o and Mistral-Large are comparably good at opinions and commonsense. For social norms and cultural commonsense, where native humans show high confidence, Google search often returns fewer relevant answers.
5 Prompting LMs for Cultural Awareness
Figure 6: Average scores of Llama3.1-8B-Instruct on native and foreign cultures when aligned using data from native, foreign, or mixed cultures. The highest performance on native culture is achieved when mixing native and foreign samples during alignment.
Beyond improving medium-sized LMs with culture-specific human preference, we evaluate the performance of larger LMs and compare with native human familiarity and web-based sources (§5.1). To better understand these differences, we examine the role of querying language (§5.2).
5.1 Main Results
@@
We follow the evaluation setup in §4.1 and assess the larger ($>$ 25B parameters) versions of the evaluated LMs with zero-shot prompting. Table 4 shows the results. Larger models are better for both Chinese and Arabic, with GPT-4o achieving the best overall performance. The more Chinese-centric model Qwen2.5-72B excels in the knowledge about cultural entities (8.42), social norms (9.06), and literacy (8.53) for Chinese; meanwhile, GPT-4o and Mistral-Large are comparably good if not better at offering opinions (9.43 and 8.96) and answering commonsense questions (9.00 and 8.58) about Chinese culture. This is interesting and likely because cultural commonsense knowledge is often unstated in the native language, while more balanced and thoughtful opinions may be expressed in foreign languages. To illustrate and test this hypothesis, we conduct further analysis in §5.2.
@@
To further examine the coverage of cultural knowledge in real-world text, we extract the main content from the top-10 ranked webpages returned by Google [^1] Google programmable search engine API: https://programmablesearchengine.google.comand evaluate their quality using the same LM-as-a-judge assessment framework, recording the highest score for each question in the CARE. Results are shown in Table 4. Since retrieved sentences from relevant web pages often fail to answer the questions directly, their scores are generally lower than LMs’ scores.

### v4

@@ 4.1 Experiment Setup
Table 4: Average scores (1: poor $\rightarrow$ 10: excellent) on Chinese, Arab, and Japanese cultures for a variety of prompting approaches, supervised fine-tuning, and preference learning using culture-specific (CARE) vs. general instruction-tuning (multilingual Alpaca) and preference (translated and filtered OpenOrca/UltraFeedback) data. SFT is performed on the instruction data only, while preference learning is conducted on the preference pairs.
Baselines.
We compare cultural preference tuning on CARE with four groups of baselines: (1) General-domain preference datasets: we adopt the multilingual translations 2A2I (2024); Pan (2024); Hwang (2024) of two synthetic English preference datasets, OpenOrca Intel (2024) (2.1k Ar, 2.0k Zh, 12.9k Ja pairs) and UltraFeedback Cui et al. (2023a) (2.0k Ar, 1.7k Zh pairs). We also compare with HelpSteer3 Wang et al. (2025b), a human-annotated multilingual preference dataset (1.1k Zh, 0.2k Ja pairs). All focus on truthfulness, honesty, and helpfulness. (2) Instruction-tuned models: we conduct SFT on our CARE and the multilingual Alpaca corpus Cui et al. (2023b); Chen et al. (2023); fujiki (2023). (3) Multilingual preference tuning: MAPO She et al. (2024). (4) Prompting methods: CoT Wei et al. (2022) and role-play Kong et al. (2023). Implementation details are in Appendix B.
LM-as-a-judge Evaluation and its Reliability.
For evaluation, we adopt the LM-as-a-judge strategy Zheng et al. (2023), prompting GPT-4o to generate a rationale and assign a 1-10 score on how aligned the response is to the human reference. We validate this setup by comparing GPT-4o’s ratings with native annotators’, where we achieve a high Pearson correlation of 0.93. More details and full evaluation prompts can be found in Appendix C.
4.2 Main Results
Table 3 presents the results of medium-sized LMs before and after cultural preference learning with CARE, while Table 4 compares results with various baselines. We see the following key findings:
Using CARE, culturally-aligned LMs achieve higher average scores (up to 29% improvement) compared to the vanilla checkpoints across all three cultures. This resonates with the findings of Zhou et al. (2024a) and shows that a relatively small amount of carefully curated data by humans can improve LMs’ alignment. We also see a noticeable gap among LMs developed by different regions. The Qwen2.5-7B developed by China-based Alibaba performs the best in Chinese and can be further improved from 7.28 to 7.61 (out of 10) by conducting preference optimization with CARE. Interestingly, the aligned version of Gemma2-9B outperforms Qwen2.5-7B on social norms and commonsense for Chinese (more on this in §5.2).
Cultural preference learning strengthens stronger LMs but fails to address weaknesses in weaker LMs. Qwen2.5-7B and Gemma2-9B, the top-performing LMs on Chinese and Arabic, respectively, show consistent improvement across cultural categories after preference learning. However, preference learning shows limited gains where the model’s initial performance is poor. Mistral-7B is not improved on Chinese entities and literacy, where its starting scores are only 3.03 and 2.43. Similarly, Qwen2.5-7B, does not benefit from preference learning on Arabic entities. Models also show limited progress on Japanese data, reflecting their shallow exposure to Japanese culture. Similarly, when the model possesses foundational cultural knowledge, utilizing the preference signals from the model itself via methods like MAPO rather than directly from human preferences can also lead to improvements. All these suggest that base models need basic cultural knowledge for preference learning to be effective.
Culture-specific human preference data is helpful beyond general human preference. Table 4 shows that tuning with CARE’s cultural preference data improves performance over much larger general preference datasets, even though they contain 3–17$\times$ more training pairs. Cultural preference learning with CARE also improves performance in both CoT and role-play prompting setups. While culture-specific SFT outperforms generic SFT, preference tuning yields greater gains, highlighting the value of human cultural preferences over ground-truth demonstrations alone.
4.3 Scaling of Data and Model Sizes
We examine how scaling both model size and data volume influences the performance before and after cultural preference learning. Specifically, we consider the Qwen2.5 {1.5B, 3B, 7B, 14B} and the Gemma2 {2B, 9B} series, which offer multiple model sizes. For data sizes, we align models using DPO on different proportions {0%, 25%, 50%, 75%, 100%} of Chinese cultural preference pairs within CARE. As shown in Figure 3, consistent improvements are observed across different model and data sizes. Even a relatively small amount (e.g., 25%) of data in CARE can lead to improvements, particularly for smaller-sized LMs. Scaling up the data leads to progressively better performance, highlighting the benefits of employing more cultural preference data in future work.
4.4 Cultural Generalization
We evaluate CARE-aligned models on other cultural tasks using culture-related data that is not included in CARE. Specifically, we use the short answer questions within the Blend Myung et al. (2024) benchmark and multi-choice questions within the Include Romanou et al. (2025) benchmark, both written in the native languages. From Figure 4, we observe that preference learning with culture-specific human preference can also benefit out-of-domain cultural tasks, indicating the generalization capability of CARE. Besides, we also test the impact of cultural preference learning on general NLP benchmarks in Appendix D.3, showing that cultural preference learning does not hinder models’ overall capabilities.
4.5 Multi-cultural Alignment
We analyze the impact of incorporating data about different cultures in preference learning. To ensure a fair comparison, we select an equal number of samples for each trial. Specifically, we align Llama3.1-8B-Instruct with 675 Chinese and 547 Arabic pairs sampled from three different contexts: native (Chinese or Arab) culture, foreign cultures (non-Chinese or non-Arab), and mixed pairs (half native and half foreign). Results are shown in Figure 5. Tuning on foreign-culture pairs alone lowers native performance; adding native pairs raises them (Chinese 4.69 $\rightarrow$ 4.87, Arab 2.98 $\rightarrow$ 3.50), and combining native and foreign pairs further improves performance, indicating that geographically diverse preference data can further strengthen cultural awareness.
Figure 4: Accuracy on the Chinese, Arabic, and Japanese subsets of the short answer questions in Blend Myung et al. (2024) and multi-choice questions in Include Romanou et al. (2025). CARE preference tuning improves out-of-domain cultural tasks for most cultures and models, except on the Arabic subset of Include. Blend does not provide a Japanese subset.
Figure 5: Average scores of Llama3.1-8B-Instruct on local and foreign cultures when aligned using data from native, foreign, or mixed cultures. The highest performance on local culture is achieved when mixing local and foreign samples during preference learning.
4.6 Language-specific LLMs
We also explore the combined benefits of both language-specific training and cultural preference tuning. In particular, we conduct DPO tuning on CARE’s Japanese training split with the Swallow model variants (Llama3.1-Swallow-8B, Gemma2-Swallow-9B, and Mistral-Swallow-7B), which are continuously pre-trained and instruction-tuned on the Japanese corpus Fujii et al. (2024). Table 6 shows that language-specific continual pre-training improves each base model by 1–3 points (on a 10-point scale), while subsequent preference tuning on CARE contributing an additional 0.2–0.4 points. Combined, these steps enable smaller models to outperform much larger LLMs. For example, aligned Llama-Swallow-8B outperforms vanilla Llama-70B (6.01 vs. 5.82), and aligned Gemma-Swallow-9B surpasses Gemma-27B (6.61 vs. 5.85). The results echo our findings in §4.2: cultural alignment is consistently helpful, even when base models have enough language proficiency.
@@ 4.6 Language-specific LLMs
Table 5: Average scores on CARE of larger LMs and web content, evaluated by the judge LM across all samples (1: poor $\rightarrow$ 10: excellent). “Human Awareness (%)” indicates the percentage of questions for which locals know the correct answer. Chinese-developed Deepseek-v3 and Qwen2.5 excel on Chinese entities, norms, and literacy, while GPT-4o and Mistral-Large are comparably good at opinions and commonsense. For social norms and cultural commonsense, where local humans show high confidence, Google search often returns fewer relevant answers.
Model Entities Opinion Norms C. sense Literacy Avg. Llama family Llama3.1-8B 2.20 3.93 2.47 2.57 1.97 2.63 $\hookrightarrow$ CARE Aligned[Uncaptioned image] 3.20 4.67 2.80 3.20 1.67 3.11 $\hookrightarrow$ Swallow-8B 5.63 6.40 5.47 6.33 4.63 5.69       $\hookrightarrow$ CARE Aligned[Uncaptioned image] 5.77 6.57 5.60 6.70 5.40 6.01 Llama3.3-70B 6.17 6.07 4.77 6.13 5.97 5.82 $\hookrightarrow$ Swallow-70B 7.20 7.37 7.13 7.53 7.27 7.30 Gemma family Gemma2-9B 4.46 6.50 5.67 6.50 2.33 5.09 $\hookrightarrow$ CARE Aligned[Uncaptioned image] 4.40 6.60 5.80 6.67 2.93 5.28 $\hookrightarrow$ Swallow-9B 6.07 7.27 6.90 7.20 5.10 6.51       $\hookrightarrow$ CARE Aligned[Uncaptioned image] 6.10 7.13 7.37 6.97 5.50 6.61 Gemma2-27B 5.37 7.07 6.50 6.47 3.83 5.85 $\hookrightarrow$ Swallow-27B 6.37 7.43 6.43 7.43 6.00 6.73 Mistral family Mistral-7B 1.70 3.96 2.40 2.48 1.20 2.34 $\hookrightarrow$ CARE Aligned[Uncaptioned image] 1.83 3.90 2.43 2.33 1.23 2.35 $\hookrightarrow$ Swallow-7B 2.70 4.43 4.07 3.87 2.00 3.41       $\hookrightarrow$ CARE Aligned[Uncaptioned image] 3.13 4.00 4.50 4.90 2.53 3.81 Mistral-8×7B 2.10 4.37 3.23 3.30 2.30 3.06 $\hookrightarrow$ Swallow-8×7B 4.33 5.40 5.30 5.87 2.90 4.76
Table 6: Average scores on CARE’s Japanese test set for (i) each backbone, (ii) each backbone further tuned on CARE, (iii) its Japanese-pretrained Swallow variant, and (iv) that variant further tuned on CARE.
5 Prompting LMs for Cultural Awareness
We evaluate larger LMs and compare to human familiarity and web-based sources (§5.1), as well as querying in different languages (§5.2).
@@ 5.1 Main Results
We follow the evaluation setup in §4.1 and assess the larger ($>$ 25B parameters) versions of the evaluated LMs with zero-shot prompting. Table 5 shows the results. Larger models are better for all three cultures, with closed-source ones like GPT-4o and Deepseek-v3 leading the overall performance. The more Chinese-centric model Deepseek-v3 and Qwen2.5-72B excels in the knowledge about cultural entities (8.96 and 8.42), social norms (9.03 and 9.06), and literacy (9.03 and 8.53) for Chinese; meanwhile, GPT-4o and Mistral-Large are comparably good if not better at offering opinions (9.43 and 8.96) and answering commonsense questions (9.00 and 8.58) about Chinese culture. This is interesting and likely because cultural commonsense knowledge is often unstated in the native language, while more balanced and thoughtful opinions may be expressed in foreign languages. We further test this hypothesis in §5.2.
@@
To further examine the coverage of cultural knowledge in real-world text, we extract the main content from the top-10 ranked webpages returned by Google111 https://programmablesearchengine.google.comand evaluate their quality using the same LM-as-a-judge assessment framework, recording the highest score for each question in the CARE. Results are shown in Table 5. Since retrieved sentences from relevant web pages often fail to answer the questions directly, their scores are generally lower than LMs’ scores.